# Optimizing the interaction geometry of inverse Compton scattering x-ray sources

### v1

@@ Abstract
Inverse Compton scattering is developing to be a promising method to generate coherent and tunable x-rays. In this paper we present a theoretical framework to describe an inverse Compton scattering x-ray source for arbitrary interaction angles between the electron and laser pulse. Importantly the divergence of a tightly focused laser pulse will have a significant impact of the number of scattered x-rays. The parameters of the interaction geometry are optimized for two specific cases: head-on scattering; and a grazing co-propagating interaction angle. For head-on scattering we show that a tight symmetrically focused laser pulse, that balances laser intensity and interaction time, optimizes the x-ray brilliance. For a grazing angle geometry an elliptical focus of the laser pulse is required to mitigate a reduced interaction time. We find that the latter geometry is especially useful for soft x-ray generation.
@@ I Introduction
Figure 1: a) Schematic representation of ICS. An electron (green) interacts with a laser pulse (red) under an angle $\theta_{L}$, generating an x-ray with emission angle $\theta_{\mathrm{x}}$. b) shows the angular radiation distribution, which becomes symmetrical for $\theta_{\mathrm{x}}\ll 1/\gamma$. Here the vertical axis is in the direction of the laser polarization. c) shows the spectral angular density, with $\omega_{0}=\omega_{L}(1-\beta\cos\theta_{L})/(1-\beta)$ is the on-axis x-ray frequency.
@@
Coherent and tunable x-ray sources have become a central tool in many field of science, such as materials science [sedigh2020review], life sciences [thomlinson2018k], and advanced medical diagnostics [quenot2022x]. This demand is primarily served by large facilities such as synchrotrons and free electron lasers [couprie2014new]. However, due to the limited availability of these large facilities, their usage comes with many logistical challenges. Therefore, there is a growing need for coherent and tunable x-ray sources at the scale of the university lab.
@@
Inverse Compton scattering (ICS), in which relativistic electrons collide with an intense laser pulse, forms the basis of a promising method to generate x-rays at the lab scale. Several ICS based x-ray sources are currently operational [gunther2020versatile, du2013generation, dupraz2020thomx, graves2014compactburstmode, VanElk:25, barty2024design, amoudry2025commissioning] and new designs have been proposed [faillace2019status, brummer2022compact, musat2024a, schaap2024inverse]. Imaging techniques that require energy tunability and coherence, such as K-edge subtraction imaging [gunther2025rapid] and phase contrast imaging [Berthe_2024_PhaseContrast], have been demonstrated with an ICS based x-ray source and the use of lensless techniques such as ptychography has also been proposed [batey2021x].
@@
While the fundamental theory of ICS has been described in great detail [hartemann2013nonlinear, esarey1993nonlinear, schaap2022photon], a closed form analytical treatment, that takes into account both the three dimensional distribution of the pulsed electron and laser beams and arbitrary interaction angles, is lacking. Optimization tasks are therefore left to simulations, which can be time consuming, and more importantly, do not provide a thorough physical understanding.
@@
The goal of this paper is to provide an analytical closed form theoretical framework to optimize x-ray generation by ICS. The framework is meant to provide intuitive understanding on how the x-ray source performance scales with the parameters of the interaction geometry. Here, the term geometry refers not only the interaction angle, but also the electron and laser pulse dimensions, the electron beam energy and the bunch charge.
The framework could serve as a baseline for a Compton x-ray source design, where given a desired x-ray energy and performance of the electron beamline and laser system, the optimized interaction geometry can be calculated directly.
I.1 Basic properties of inverse Compton scattering
@@
Figure 1a illustrates the geometry of ICS. An electron (green) with velocity $\vec{v}=\beta c\rm\hat{z}$ and normalized energy $\gamma=1/\sqrt{1-\beta^{2}}$ interacts with a laser pulse (red) under an interaction angle $\theta_{L}$, w.r.t. the electron velocity. This angle is defined such that $\theta_{L}=\pi$ corresponds to a head-on collision. During the interaction, the electron starts to rapidly oscillate, radiating in the process (purple). Due to the relativistic nature of the electron, the radiation becomes strongly blue shifted. The frequency up-shift is given by [schaap2022photon]
@@
Here, $\omega_{\mathrm{x}}$ and $\omega_{L}$ are the angular frequencies of the radiation and the laser pulse, respectively, and $\theta_{\mathrm{x}}$ is the angle of the outgoing radiation with respect to the electron velocity.
In the relativistic limit ($\gamma\gg 1$) and under small radiation angles ($\gamma\theta_{\mathrm{x}}\ll 1$), Eq. (1) can be approximated. For head-on scattering $(\theta_{L}=\pi)$ this results in the well known expression
@@
Note here that, since typically a large frequency up-shift is desired, $\gamma\theta_{\mathrm{L}}\gg 1$. These approximations clearly show that the x-ray energy can be tuned by varying either the electron energy, $\gamma$, or the interaction angle $\theta_{L}$.
Additionally, Eqs. 2 and 3 show how the x-ray frequency depends on the angle $\theta_{\mathrm{x}}$, with the highest frequency emitted along the propagation direction of the electron ($\theta_{\mathrm{x}}=0$) and the frequency is progressively decreasing for larger angles, as shown in Fig 1c. At $\theta_{\mathrm{x}}\simeq 1/\gamma$ the x-ray frequency has become half of the on-axis frequency. Fig 1b shows the intensity distribution of ICS, which is concentrated in a narrow cone characterized by a half angle $\theta_{\mathrm{x}}=1/\gamma$ [jackson_classical_1999]. The distribution is slightly asymmetric for linear laser polarization, which is in Fig 1b in the vertical direction. However, for $\theta_{\mathrm{x}}\ll 1/\gamma$, which is required for a narrow bandwidth x-ray beam, the radiation distribution is to a good approximation uniform.
@@
with $r_{e}$ the classical electron radius. Obviously, for efficient generation of x-rays with a Compton source the density of electrons and laser photons needs to be as high as possible, requiring focused electron and laser pulses.
@@ I.2 Structure of this paper
We will analyze the dependence of the x-ray generation on geometry of ICS in increasing levels of complexity, illustrated by Fig. 2. We start in Sec. II by describing ICS in a head-on geometry using a collisional model. We first consider the collision between pencil beams, i.e. in the approximation of perfectly parallel beams (1 in Fig. 2). Next, we analyze the more realistic situation of a focused laser pulse and a pencil electron beam (2 in Fig. 2) and find the optimized interaction spot size and a corresponding optimized bunch charge.
In Sec. III, we analyze ICS under arbitrary interaction angles. This is the most novel part of the paper. A proper analysis of this geometry requires a treatment in a covariant framework and reveals in several subtle effects. Again, we the situation of two pencil beams (3 in Fig. 2) subsequently the situation of a focused laser pulse and a pencil electron beam (4 in Fig. 2). We derive closed form analytical expressions for the parameters that define interaction geometry which optimize the x-ray brilliance in a grazing angle geometry. We find that a shallow co-propagating interaction angle – a grazing angle geometry – can be favorable compared to a head-on geometry, especially for generation of soft x-rays.
Figure 2: Four illustrations of the geometries considered throughout this paper. In each geometry the electron bunch (green) with velocity $\bm{\beta}$ and laser pulse (red) with wave vector $\bm{k}_{L}$ are shown at three different times up until the collision. The top figures show the pencil beam (1) and focused laser beam (2) description in a head-on geometry. The bottom figures the pencil beam (3) and focused laser beam (4) description in a grazing angle geometry. The relevant r.m.s. pulse dimensions are also indicated
@@ II Head-on inverse Compton scattering
In this section we describe ICS in a head-on geometry $(\theta_{L}=\pi)$. We start with a pencil beam description and calculate the x-ray brilliance. The result from this simplified model serves as a demonstration of important scalings for ICS. Next, we consider a finite laser pulse length to find the optimal laser waist. Finally, a simple argument is used to find the bunch charge that maximizes the x-ray brilliance.
@@ II.1 Pencil beam description
The total number of scattered photons from ICS can be calculated by considering an electron pulse and a laser pulse colliding in a head-on geometry. First we assume the pulses to be pencil beams with a Gaussian transverse photon or electron density
$\rho_{L,e}=\frac{N_{L,e}}{2\pi\sigma^{2}_{L,e}}\exp\left(-\frac{x^{2}+y^{2}}{2\sigma_{L,e}^{2}}\right),$ (5)
where the subscripts $L$ and $e$ refer to either the laser or electron beam. $N_{L,e}$ is the total number of laser photons or electrons and $\sigma_{L,e}$ is the r.m.s. transverse size. Throughout this paper a quantity $\sigma$ refers to an r.m.s. quantity (with the exception of the Thompson cross section $\sigma_{T}$). Note that we can neglect the longitudinal distributions due to the head-on geometry and the pencil beam approximation. The total number of scattered x-rays in the collision is given by [furman2003moller]
@@
Note that Eq. (6) considers the x-rays scattered in the complete $4\pi$ solid angle and the corresponding full bandwidth. If a small relative energy spread $\Delta\omega/\omega_{\mathrm{x}}$ is required, the x-ray acceptance angle should be reduced to $\Theta_{\mathrm{x}}=\sqrt{(\Delta\omega/\omega_{\mathrm{x}})}/\gamma$. If $\gamma\Theta_{\mathrm{x}}\ll 1$, the number of x-rays in this acceptance angle is [jackson_classical_1999]
@@
Here, we have additionally assumed the size of the electron and laser pulses to be matched ($\sigma_{L}=\sigma_{\mathrm{e}}$), which is an important assumption throughout this section. A figure of merit often used to compare x-ray sources is the x-ray brilliance, which takes into account the photon flux, transverse coherence and relative energy spread. The average r.m.s. x-ray brilliance is defined as [attwood2000soft]
Here, $f$ is the repetition rate, $\Delta A$ is the x-ray source area and the $\Delta\Omega$ is the solid angle in which the x-rays are emitted. For details regarding the definition of the x-ray brilliance, see Appendix C. Using $\Delta A\Delta\Omega=\pi^{2}\sigma_{\mathrm{e}}^{2}\Theta_{\mathrm{x}}^{2}$ the average r.m.s brilliance becomes
$B_{\mathrm{x}}=\frac{f\sigma_{\mathrm{T}}N_{L}N_{\mathrm{e}}}{4\pi^{3}}\frac{3}{2}\frac{1}{\Theta_{\mathrm{x}}^{2}\sigma_{e}^{4}}\frac{\gamma^{2}\Theta_{\mathrm{x}}^{2}}{\Delta\omega/\omega_{\mathrm{x}}}.$ (9)
The x-ray energy spread of a Compton x-ray source has several contributions originating from both the laser pulse and electron beam [petrillo2012photon]. For now, to illustrate important scalings, we make a simplification and consider only the bandwidth due to the angular spread of the electron beam $\left.\Delta{\omega}/\omega_{\mathrm{x}}\right|_{\sigma_{\theta\mathrm{x}}}=\gamma^{2}\sigma_{\theta e}^{2}$. If we describe the electron beam using the normalized emittance, $\epsilon_{n}=\gamma\beta\sigma_{e}\sigma_{\theta e}\simeq\gamma\sigma_{e}\sigma_{\theta e}$ , Eq. (9) can be rewritten as
$B_{\mathrm{x}}=\frac{f\sigma_{\mathrm{T}}N_{L}N_{\mathrm{e}}}{4\pi^{3}}\frac{3}{2}\frac{\gamma^{4}\sigma_{\theta e}^{2}}{\epsilon_{n}^{4}}.$ (10)
Equation (10) shows two important scalings: First, we find that the brilliance scales with $\gamma^{4}$, which originates from the improved geometric emittance with increased electron energy. This not only results in a smaller x-ray source size, but also allows for tighter focusing of the laser pulse. In a head-on geometry, the electron energy is fixed by the desired x-ray energy and laser wavelength. If the interaction angle is varied, however, this scaling can be exploited for any desired x-ray energy. This will be further explored in Sec. III.
Second, Eq. (10) shows that $B_{\mathrm{x}}\propto N_{\mathrm{e}}/\epsilon_{n}^{4}$. For space charge dominated electron beams, it holds that $N_{\mathrm{e}}/\epsilon_{\mathrm{n}}^{2}=\mathrm{constant.}$ As a result we find that the x-ray brilliance increases with a lower bunch charge. This counterintuitive result is a direct consequence of the assumption that the spot size of the electron and laser beams are matched $(\sigma_{L}=\sigma_{e})$, which is an assumption that cannot hold indefinitely, as it is obviously limited by the diffraction limit. In the remainder of this section we calculate the x-ray flux and brilliance with realistic electron and laser beam descriptions, and find the optimized laser focus and the corresponding bunch charge to match the electron spot size to the laser pulse.

### v2

@@ Abstract
Inverse Compton scattering (ICS) is a promising method for generating coherent and tunable x-rays in a compact setup. In this paper, we present a theoretical framework describing the output of an ICS x-ray source for arbitrary interaction angles between pulsed electron and laser beams. This allows for analytic optimization of the x-ray beam properties by varying the parameters defining the geometry. In general, different x-ray applications require optimization of different x-ray beam properties, such as energy spread for x-ray spectroscopy and angular spread for x-ray scattering measurements. In this paper, we restrict ourselves to optimization of the x-ray brilliance, which is a comprehensive figure of merit for x-ray beam quality. We investigate two specific ICS interaction geometries in particular: head-on scattering of a laser beam off an electron beam; and scattering of a laser beam off an electron beam in a co-propagating geometry, interacting under a grazing angle. For head-on scattering we show that a tightly focused, cylindrically symmetric laser pulse, which balances laser intensity and interaction time, optimizes the x-ray brilliance. For a co-propagating, grazing angle geometry, an elliptical focus of the laser pulse is required to mitigate the effect of a reduced interaction time. We find that the latter geometry is especially useful for soft x-ray generation.
@@ I Introduction
Figure 1: a) Schematic representation of ICS. An electron (green) interacts with a laser pulse (red) under an angle $\theta_{L}$, generating x-rays with emission angle $\theta_{\mathrm{x}}$. b) shows the angular radiation distribution, which becomes cylindrically symmetric for $\theta_{\mathrm{x}}\ll 1/\gamma$. Here the vertical axis is in the direction of the laser polarization. c) shows the spectral angular density, with $\omega_{0}=\omega_{L}(1-\beta\cos\theta_{L})/(1-\beta)$ the x-ray frequency for $\theta_{\mathrm{x}}=0$
@@
Coherent and tunable x-ray sources have become an indispensable tool in many fields of research, such as materials science [38], life sciences [42], and advanced medical diagnostics [33]. This demand is primarily served by large facilities such as synchrotrons and free electron lasers [12]. However, due to the limited availability of these large facilities, their usage comes with many logistical challenges. Therefore, there is a growing need for coherent and tunable x-ray sources at the scale of the university lab.
@@
Inverse Compton scattering (ICS), a phenomenon in which a highly energetic electron collides with a photon in such a way that the electron looses energy and the photon gains energy, forms the basis of a promising method to generate x-rays at the lab scale. Several ICS based x-ray sources are currently operational [19, 13, 3, 18, 14, 6, 4] and new designs have been proposed [16, 10, 27, 35]. Imaging techniques that require energy tunability and coherence, such as K-edge subtraction imaging [20] and phase contrast imaging [9], have been demonstrated with an ICS based x-ray source and the use of lensless techniques such as ptychography has been proposed [7].
@@
While the fundamental theory of ICS has been described in great detail [21, 15, 37], a closed form analytical treatment, that takes into account both the three dimensional distribution of the pulsed electron and laser beams and arbitrary interaction angles, is lacking. Optimization tasks are therefore left to simulations, which can be time consuming and do not provide a thorough physical understanding of the limitations.
The goal of this paper is to provide an analytical closed form theoretical framework to optimize x-ray generation by ICS. Different applications require optimization of different x-ray beam properties. In this paper we choose to optimize the x-ray brilliance, which is a comprehensive figure of merit that take into account x-ray flux, transverse coherence and energy spread. The average x-ray brilliance is defined as [5]
Here, $N_{\mathrm{x}}$ is the number of x-rays per pulse, $f$ is the pulse repetition frequency, $\Delta A$ is the effective x-ray emission area, and $\Delta\Omega$ is the solid angle in which the x-rays are emitted. $\Delta\omega/\omega_{\mathrm{x}}$ is the relative bandwidth of the x-ray beam. For details on the definition of the x-ray brilliance used in the paper, see Appendix C.
@@
The framework is meant to provide an intuitive understanding of how the x-ray source performance scales with the parameters of the interaction geometry. Here, the term geometry refers not only to the interaction angle, but also to the electron and laser pulse dimensions, the electron beam energy, and the bunch charge. The framework could serve as a baseline for a Compton x-ray source design, in which, given a desired x-ray energy and performance of the electron beamline and laser system, the optimized interaction geometry can be calculated directly.
I.1 Basics of inverse Compton scattering
@@
Figure 1a shows a schematic illustration of the ICS geometry. An electron (green) with velocity $\vec{v}=\beta c\rm\hat{z}$ and normalized energy $\gamma=1/\sqrt{1-\beta^{2}}$ interacts with a laser pulse (red) under an interaction angle $\theta_{L}$, w.r.t. the electron velocity. This angle is defined in such a way that $\theta_{L}=\pi$ corresponds to a head-on collision. During the interaction, the electron oscillates rapidly and radiates in the process (purple). Due to the relativistic nature of the electron, the radiation becomes strongly blue shifted. The shifted frequency is given by [37]
@@
Here, $\omega_{\mathrm{x}}$ and $\omega_{L}$ are the angular frequencies of the generated radiation and the laser pulse, respectively, and $\theta_{\mathrm{x}}$ is the angle of the outgoing radiation with respect to the electron velocity.
In the relativistic limit ($\gamma\gg 1$) and under small radiation angles ($\gamma\theta_{\mathrm{x}}\ll 1$), Eq. (2) can be approximated in a convenient way. For head-on scattering $(\theta_{L}=\pi)$ this results in the well known expression
@@
Note that, since typically a large frequency shift is desired, preferably $\gamma\theta_{\mathrm{L}}\gg 1$. These approximations clearly show that the x-ray energy can be tuned by varying either the electron energy, $\gamma$, or the interaction angle $\theta_{L}$.
Furthermore, Eqs. 3 and 4 show how the x-ray frequency depends on the angle $\theta_{\mathrm{x}}$: the highest x-ray frequency, $\omega_{0}$, is emitted along the propagation direction of the electron ($\theta_{\mathrm{x}}=0$) and the x-ray frequency gradually decreases for larger angles, as shown in Fig 1c. At $\theta_{\mathrm{x}}\simeq 1/\gamma$ the x-ray frequency has become half of the on-axis frequency, $\omega_{0}$. Fig 1b shows the intensity distribution of ICS, which is concentrated in a narrow cone characterized by a half angle $\theta_{\mathrm{x}}=1/\gamma$ [23]. The distribution is slightly asymmetric for linear laser polarization, which is pointing in the vertical direction in Fig 1b. However, for $\theta_{\mathrm{x}}\ll 1/\gamma$, which is required for a narrow bandwidth x-ray beam, the angular distribution of the radiation is to a good approximation uniform.
@@
where $r_{e}=e^{2}/(4\pi\epsilon_{0}mc^{2})$ is the classical electron radius, with $e$ the elementary charge and $m$ the electron mass. Obviously, for efficient generation of x-rays by Compton scattering, the density of electrons and laser photons needs to be as high as possible, requiring focused electron and laser pulses.
@@ I.2 Structure of this paper
We will analyze the dependence of the x-ray generation on geometry of ICS in increasing levels of complexity, illustrated by Fig. 2. We start in Sec. II by describing ICS in a head-on geometry using a collisional model. We first consider the collision between pencil beams, i.e. in the approximation of perfectly parallel beams (1 in Fig. 2). Next, we analyze the more realistic situation of a focused laser pulse and a pencil electron beam (2 in Fig. 2) and find the optimized interaction spot size and the corresponding optimized bunch charge.
In Sec. III, we analyze ICS for arbitrary interaction angles. This is the most novel part of the paper. A proper analysis of this geometry requires a treatment in a covariant framework. Again, we first analyze the situation of two pencil beams (3 in Fig. 2). Subsequently, we investigate the combination of a focused laser pulse and a pencil electron beam (4 in Fig. 2). We derive closed form analytical expressions for the laser focal spot size and the electron bunch charge, which optimize the x-ray brilliance in a grazing angle geometry. We find that a shallow co-propagating interaction angle – a grazing angle geometry – can be favorable compared to a head-on geometry, especially for the generation of soft x-rays.
Figure 2: Schematic illustrations of the four geometries considered in this paper. In each geometry the electron bunch (green) with velocity $\bm{\beta}$ and laser pulse (red) with wave vector $\bm{k}_{L}$ are shown at three different times up until the collision. The top figures show the pencil beam 1 and focused laser beam 2 geometries in a head-on geometry. The bottom figures show the pencil beam 3 and focused laser beam 4 geometries in a grazing angle geometry. The relevant r.m.s. pulse dimensions are also indicated
@@ II Head-on inverse Compton scattering
In this section we describe ICS in a head-on geometry $(\theta_{L}=\pi)$. We start with a pencil beam description and calculate the x-ray brilliance. The result from this simplified model serves as a demonstration of important scalings for ICS. Next, we consider a finite laser pulse length to find the optimal laser waist. Finally, a simple argument is used to determine the bunch charge that maximizes the x-ray brilliance.
@@ II.1 Pencil beam description
We assume the pulses to be pencil beams with a Gaussian transverse photon/electron density distribution
$\rho_{L,e}=\frac{N_{L,e}}{2\pi\sigma^{2}_{L/e}}\exp\left(-\frac{x^{2}+y^{2}}{2\sigma_{L/e}^{2}}\right),$ (6)
where the subscripts $L$ and $e$ refer to the laser beam and the electron beam, respectively. $N_{L/e}$ is the total number of laser photons/electrons and $\sigma_{L/e}$ is the r.m.s. transverse size. Throughout this paper, a quantity $\sigma$ refers to an r.m.s. quantity (with the exception of the Thompson cross section $\sigma_{T}$). Note that we can neglect the $z$-dependence of the density distributions due to the head-on geometry and the pencil beam approximation. The total number of scattered x-rays in the collision is given by [17]
@@
Note that Eq. (7) corresponds to the situation in which the x-rays are scattered in the complete $4\pi$ solid angle with 100% relative energy spread. If a small relative energy spread $\Delta\omega/\omega_{\mathrm{x}}$ is required, the x-ray acceptance angle should be reduced to $\Theta_{\mathrm{x}}=\sqrt{(\Delta\omega/\omega_{\mathrm{x}})}/\gamma$. If $\gamma\Theta_{\mathrm{x}}\ll 1$, the number of x-rays in this acceptance angle is [23]
@@
Here, we have additionally assumed that the sizes of the electron and laser pulses are matched ($\sigma_{L}=\sigma_{\mathrm{e}}$), which is close to the optimized situation and simplifies the analysis. This is an important assumption used throughout this section. Using $\Delta A\Delta\Omega=2\pi^{2}\sigma_{\mathrm{e}}^{2}\Theta_{\mathrm{x}}^{2}$ and $\Delta\omega/\omega_{\mathrm{x}}=\sqrt{2\pi}\sigma_{\omega}/\omega_{\mathrm{x}}$, with $\sigma_{\omega}/\omega_{\mathrm{x}}$ the r.m.s. relative bandwidth, the average brilliance is given by
$B_{\mathrm{x}}=\frac{f\sigma_{\mathrm{T}}N_{L}N_{\mathrm{e}}}{\sqrt{2\pi}\ 8\pi^{3}}\frac{3}{2}\frac{1}{\sigma_{e}^{4}\Theta_{\mathrm{x}}^{2}}\frac{\gamma^{2}\Theta_{\mathrm{x}}^{2}}{\sigma_{\omega}/\omega_{\mathrm{x}}}.$ (9)
The energy spread of a Compton x-ray source has several contributions originating from both the laser pulse and the electron beam properties [29]. To illustrate important scalings, we restrict ourselves in the first instance to the contribution to the energy spread due to the angular spread of the electron beam: $\left.\sigma_{\omega}/\omega_{\mathrm{x}}\right|_{\sigma_{\theta e}}=\gamma^{2}\sigma_{\theta e}^{2}$. If we describe the electron beam using the normalized emittance, $\epsilon_{n}=\gamma\beta\sigma_{e}\sigma_{\theta e}\simeq\gamma\sigma_{e}\sigma_{\theta e}$ , Eq. (9) can be rewritten as
$B_{\mathrm{x}}=\frac{f\sigma_{\mathrm{T}}N_{L}N_{\mathrm{e}}}{\sqrt{2\pi}\ 8\pi^{3}}\frac{3}{2}\frac{\gamma^{4}\sigma_{\theta e}^{2}}{\epsilon_{n}^{4}}.$ (10)
Equation (10) shows two important scalings: First, we find that the brilliance scales with $\gamma^{4}$, which originates from the improved geometric emittance with increased electron energy. This not only results in a smaller x-ray source size, but also allows for tighter focusing of the laser pulse. In a head-on geometry, the electron energy is fixed by the desired x-ray energy for a given laser wavelength. However, if the interaction angle can be varied, this scaling can be exploited for any desired x-ray energy. This will be further explored in Sec. III.
Second, Eq. (10) shows that $B_{\mathrm{x}}\propto N_{\mathrm{e}}/\epsilon_{n}^{4}$. For space charge dominated electron beams, it holds that $N_{\mathrm{e}}/\epsilon_{\mathrm{n}}^{2}=\mathrm{constant.}$ As a result we find that the x-ray brilliance increases with a lower bunch charge. This counterintuitive result is a direct consequence of the assumption that the spot size of the electron and laser beams are matched $(\sigma_{L}=\sigma_{e})$, which is an assumption that cannot hold indefinitely, as it is obviously limited by the laser beam diffraction limit. In the remainder of this section we calculate the x-ray flux and brilliance with realistic electron and laser beam descriptions, and find the optimized laser focus and the corresponding bunch charge to match the electron spot size to the laser pulse.